\begin{thebibliography}{10}

\bibitem{abatzoglou1979}
Theagenis~J. Abatzoglou, \emph{Norm derivatives on spaces of operators}, Math.
  Ann. \textbf{239} (1979), no.~2, 129--135. \MR{519008}

\bibitem{bhatia--semrl1999}
Rajendra Bhatia and Peter \v Semrl, \emph{Orthogonality of matrices and some
  distance problems}, vol. 287, 1999, Special issue celebrating the 60th
  birthday of Ludwig Elsner, pp.~77--85. \MR{1662861}

\bibitem{Davidson-book-nestalgebras}
Kenneth~R. Davidson, \emph{Nest algebras}, Pitman Research Notes in Mathematics
  Series, vol. 191, Longman Scientific \& Technical, Harlow; copublished in the
  United States with John Wiley \& Sons, Inc., New York, 1988, Triangular forms
  for operator algebras on Hilbert space. \MR{972978}

\bibitem{Effros--Ruan-book}
Edward~G. Effros and Zhong-Jin Ruan, \emph{Operator spaces}, London
  Mathematical Society Monographs. New Series, vol.~23, The Clarendon Press
  Oxford University Press, New York, 2000. \MR{MR1793753 (2002a:46082)}

\bibitem{farenick--plosker--smith2011}
Douglas Farenick, Sarah Plosker, and Jerrod Smith, \emph{Classical and
  nonclassical randomness in quantum measurements}, J. Math. Phys. \textbf{52}
  (2011), no.~12, 122204, 26. \MR{2907638}

\bibitem{james1947}
Robert~C. James, \emph{Orthogonality and linear functionals in normed linear
  spaces}, Trans. Amer. Math. Soc. \textbf{61} (1947), 265--292. \MR{21241}

\bibitem{keckic2004}
D.~J. Ke$\check{\text c}$ki$\acute{\text c}$, \emph{Orthogonality in {${\mathfrak
  S}_1$} and {${\mathfrak S}_\infty$} spaces and normal derivations}, J. Operator
  Theory \textbf{51} (2004), no.~1, 89--104. \MR{2055806}

\bibitem{keckic2005}
D.~J. Ke$\check{\text c}$ki$\acute{\text c}$, \emph{Gateaux derivative of {$B(H)$} norm}, Proc. Amer.
  Math. Soc. \textbf{133} (2005), no.~7, 2061--2067. \MR{2137872}

\bibitem{keckic2012}
D.~J. Ke$\check{\text c}$ki$\acute{\text c}$, \emph{Orthogonality and smooth points in {$C(K)$} and
  {$C_b(\Omega)$}}, Eurasian Math. J. \textbf{3} (2012), no.~4, 44--52.
  \MR{3040686}

\bibitem{kittaneh1996}
Fuad Kittaneh, \emph{Operators that are orthogonal to the range of a
  derivation}, J. Math. Anal. Appl. \textbf{203} (1996), no.~3, 868--873.
  \MR{1417136}

\bibitem{lumer1961}
G.~Lumer, \emph{Semi-inner-product spaces}, Trans. Amer. Math. Soc.
  \textbf{100} (1961), 29--43. \MR{133024}

\bibitem{magajna1993}
Bojan Magajna, \emph{On the distance to finite-dimensional subspaces in
  operator algebras}, J. London Math. Soc. (2) \textbf{47} (1993), no.~3,
  516--532. \MR{1214913}

\bibitem{maher1992}
P.~J. Maher, \emph{Commutator approximants}, Proc. Amer. Math. Soc.
  \textbf{115} (1992), no.~4, 995--1000. \MR{1086335}

\bibitem{narcowich--ward1982}
Francis~J. Narcowich and Joseph~D. Ward, \emph{Support functions for matrix
  ranges: analogues of {L}umer's formula}, J. Operator Theory \textbf{7}
  (1982), no.~1, 25--49. \MR{650191}

\bibitem{Paulsen-book}
Vern Paulsen, \emph{Completely bounded maps and operator algebras}, Cambridge
  Studies in Advanced Mathematics, vol.~78, Cambridge University Press,
  Cambridge, 2002. \MR{MR1976867 (2004c:46118)}

\bibitem{Rockafellar-book}
R.~Tyrrell Rockafellar, \emph{Convex analysis}, Princeton Mathematical Series,
  No. 28, Princeton University Press, Princeton, N.J., 1970. \MR{0274683}

\bibitem{Singer-book}
Ivan Singer, \emph{Best approximation in normed linear spaces by elements of
  linear subspaces}, Die Grundlehren der mathematischen Wissenschaften, vol.
  Band 171, Publishing House of the Academy of the Socialist Republic of
  Romania, Bucharest; Springer-Verlag, New York-Berlin, 1970, Translated from
  the Romanian by Radu Georgescu. \MR{270044}

\bibitem{singla2021}
Sushil Singla, \emph{Gateaux derivative of {$C^\ast$} norm}, Linear Algebra
  Appl. \textbf{629} (2021), 208--218. \MR{4295982}

\bibitem{stampfli1970}
Joseph~G. Stampfli, \emph{The norm of a derivation}, Pacific J. Math.
  \textbf{33} (1970), 737--747. \MR{265952}

\bibitem{winkler1996}
Soren Winkler, \emph{Matrix convexity}, ProQuest LLC, Ann Arbor, MI, 1996,
  Thesis (Ph.D.)--University of California, Los Angeles. \MR{2694915}

\end{thebibliography}

\documentclass{amsart}

\usepackage{lmodern}
\usepackage{amsmath,amsfonts,amssymb, amsthm, xcolor, enumitem, graphicx, hyperref}
\usepackage{url}
\usepackage{appendix}

\newtheorem{theorem}[subsection]{Theorem}

\newtheorem{corollary}[subsection]{Corollary}

\theoremstyle{definition} 

\newtheorem{remark}[subsection]{Remark}

\newcommand{\bra}{\langle}
\newcommand{\ket}{\rangle}
\newcommand{\tens}[1]{\mathbin{\mathop{\otimes}\limits_{#1}}}

\begin{document}


\title[Gateaux derivative in operator spaces]{Gateaux derivative of matrix norms in operator spaces and operator systems}
\author{Sushil Singla}
\address{Department of Mathematics and Statistics, University of Regina, Regina, Saskatchewan S4S 0A2, Canada}
\curraddr{}
\email{sushil@uregina.ca}
\thanks{Supported in part by the PIMS Postdoctoral Fellowship Program and the Faculty of Science, University of Regina.}

\subjclass[2020]{Primary 46G05, 46L07, 46L30; Secondary 46B20}
\keywords{Gateaux derivatives; support mapping; completely positive maps; quantum probability measure; Birkhoff-James orthogonality}

\begin{abstract}
	We find expressions for the Gateaux derivative of the matrix norms in operator spaces, and operator systems. Some applications of the results to quantum probability measures, states on C$^*$-algebras, and Birkhoff-James orthogonality are also presented.
\end{abstract}

\maketitle


\section{Introduction}

If $(\mathcal X, \|\cdot\|)$ is a normed vector space, then the norm function $x\mapsto \|x\|$ is
sublinear, and in particular, convex. Thus, for all $x,y\in \mathcal X$,
the one-sided limit
\[
\lim_{t\rightarrow 0^+}\dfrac{\|x+ty\|-\|x\|}{t}
\]
exists, 
and the resulting quantity is
known as the \emph{Gateaux derivative} of the norm function $\|\cdot\|$ at $x$ in the direction of $y$. 
One source of interest in the Gateaux derivative is its role in Banach space geometry, where the Birkhoff-James orthogonality of vector $x$ to the vector $y$
admits descriptions involving the Gateaux derivative of the norm at $x$ in the direction of $y$, see \cite[Theorem 1.4]{keckic2004}. The notion of Gateaux derivative has been used in literature for commutator approximants (see \cite{kittaneh1996, maher1992}), and in results related with the smooth points and extreme points of the norm's unit ball (see \cite{abatzoglou1979}).

Previous works have considered the Gateaux derivative of the norm function for C$^*$-norms \cite{singla2021} and the Schatten $p$-norms \cite{keckic2004}.
In the present paper, we consider the matrix norms $\|\cdot\|_n$ that arise from an operator space or an operator system.

A key idea is the notion of support mapping on an operator space $(V, \|\cdot\|_n)$. Let $M_n$ denote the space of $n\times n$ complex matrices. For $A\in M_n$, let $\Re(A):=(A+A^*)/2$
and 
$\rho^* : M_n\rightarrow\mathbb R$ be the function 
defined by 
\[
\rho^*(A)=\inf\{t\geq 0 \,\,|\,\, \Re(A)\leq tI_n\},  
\]
where $I_n\in M_n$ is the identity matrix.  If $v\in M_n(V)$, the normed vector space of matrices with entries from $V$,
then a linear map 
$\phi : V\rightarrow M_n$ is said to be a \emph{support mapping for $v$} if, for all $k\in\mathbb N$ and $w\in M_k(V)$, 
we have 
\[
\rho^*(\phi_k(w))\leq \|w\|_k
\]
and 
\[
\rho^*(\phi_n(v))=\|v\|_n,
\]
where, for any $\ell\in\mathbb N$,
 $\phi_\ell$ denotes the linear map $M_\ell(V)\rightarrow M_\ell(M_n)$ given by $[v_{ij}]_{i,j=1}^{\ell}\mapsto [\phi(v_{ij})]_{i,j=1}^{\ell}$ for all $v_{ij}\in V$.
The existence of support mappings on $(V, \|\cdot\|_n)$, for every $v\in M_n(V)$ and every $n\in\mathbb N$, 
was established by Winkler in \cite[Theorem 3.3.2]{winkler1996}. 

For $A\in M_n$, we have $\Re(A)\leq tI_n$ if and only if $\langle\Re(A)\eta, \eta\rangle\leq t$ for all unit vectors $\eta\in\mathbb C^n$. Thus, we have \[
\rho^*(A)=\sup\{\langle\Re(A)\eta, \eta\rangle \,\,|\,\, \eta\in\mathbb C^n, \|\eta\|=1\}.\]
The first goal in this paper is to express the Gateaux derivative of $\|\cdot\|_n$ in the manner presented in Theorem \ref{thm1} below, where, 
for $A\in M_n$, 
\[
M_0(A)=\{\eta\in \mathbb C^n \,\, |\,\, \|\eta\|=1\text{ and }\bra \Re(A)\eta, \eta\ket=\|A\|\}.
\]
Note that $M_0(A)$ maybe empty set, but if $A=A^*$ then $$M_0(A)=\{\eta\in \mathbb C^n \,\,:\,\, \|\eta\|=1\text{ and }\|\Re(A)\eta\|=\|A\|\}$$ is the unit ball of the subspace determined by the kernel of the matrix $A^*A-\|A\|^2I_n$.
\begin{theorem}\label{thm1}
	If $(V, \|\cdot\|_n)$ is an operator space, and 
	$v,w\in M_n(V)$. Then there exists a support mapping for $v$ such that $M_0(\phi_n(v))\neq \emptyset$ and
	 \begin{align*}\lim_{t\rightarrow 0^+}\dfrac{\|v+tw\|_n-\|v\|_n}{t}=\max\{\bra\Re(\phi_n(w))\eta, \eta\ket \,\, |\,\,  &\phi : V\rightarrow M_n \text{ is a support mapping }\\
		&\text{for } v\text{ and }\ \eta\in M_0(\phi_n(v))\}.
		\end{align*} 
\end{theorem}

 If $\mathcal S$ be an operator system with Archimedean order unit $e$, then a \emph{matrix state} on $\mathcal S$ 
is a unital completely positive map $\phi :\mathcal S\rightarrow M_k$ such that $\phi(e)=I_k$. 
For every $k\in\mathbb N$, let $S_k(\mathcal S)$ denote the set of matrix states $\phi :\mathcal S\rightarrow M_k$. For $s\in\mathcal S$, the matrix range of $s$ is the union of the sets 
\[
W_k(s)=\{A\in M_k \,\,:\,\, A=\phi(s)\ \text{ for some }\ \phi\in S_k(\mathcal S)\}.
\]
If $\mathcal S$ is the space of bounded linear operators on a Hilbert space with identity operator $I$, then for any bounded linear operator $T$ and $n=1$, $W_1(T)$ is the topological closure of the classical numerical range of $T$. It was proved by Lumer \cite{lumer1961} that
\[
\sup\{\mathrm{Re}(\lambda)\,\,|\,\, \lambda\in W_1(T)\}=\lim\limits_{t\rightarrow 0^+}\dfrac{\|I+tT\|-1}{t},
\] where $\mathrm{Re}(a+ib)=a$ for all $a,b\in\mathbb R$.
This was extended to matrix ranges by Narcowich and Ward \cite{narcowich--ward1982} as follows. For a unital
C$^*$-algebra $\mathcal A$, let $\|\cdot\|_{\sigma}$ denote the function $x\mapsto \|x\|_\sigma$ on $M_n(\mathcal A)$ defined by 
\[
\|x\|_\sigma=\sup\left\{\sqrt{s_\phi(x^*x)}\,\,|\,\, \phi\in\mathcal S_n(\mathcal A) \right\},\quad \text{for}\quad x\in M_n(\mathcal A),
\]
where $s_\phi : M_n(\mathcal A)\rightarrow\mathbb C$ is the functional 
\[
s_\phi([x_{ij}]_{i,j=1}^n)=\dfrac{1}{n}\ \phi_n([x_{ij}])_{ij},
\]
the $(i,j)$-entry of the matrix $\phi_n([x_{ij}]_{i,j=1}^n)$. 
By \cite[Proposition 3.1, Corollary 3.4]{narcowich--ward1982}
the function $x\mapsto \|x\|_\sigma$ defines a norm on $M_n(\mathcal A)$, and it has the property
that, for all $x\in \mathcal A$ and $B=[b_{ij}]_{i,j=1}^n\in M_n$, 
\[
\max\{\mathrm{Re}(\mathrm{trace}(A^tB))\,\, |\,\, A\in W_n(x)\}=n\cdot\lim\limits_{t\rightarrow 0^+}\dfrac{\|e_n+t(\overline{b}_{ij}x)_{ij}\|_\sigma-1}{t}.
\]

The following result gives a slightly different extension of Lumer's formula.

\begin{theorem}\label{cor3}
	Let $\mathcal S$ be an operator system, and let $s_1, s_2\in M_n(\mathcal S)$. Then, we have 
\begin{align}\label{e:qaz}
\lim\limits_{t\rightarrow 0^+}\dfrac{\|s_1+ts_2\|_n-\|s_1\|_n}{t} = \frac{1}{\|s_1\|_n}\max\{& \mathrm{Re}(\bra\phi_n(s_2)\eta, \phi_n(s_1)\eta\ket) \,\, |\,\,  \phi \in S_n(\mathcal S),\\
&\|\phi_n(s_1)\|=\|s_1\|_n\text{ and } \eta\in M_0(\phi_n(s_1))\}\nonumber.
\end{align}
\end{theorem}

Let $e_n\in M_n(\mathcal S)$ denotes the diagonal matrix with $e$ at each diagonal entry. Then for $s\in M_n(\mathcal S)$, the above theorem reads as \[\lim\limits_{t\rightarrow 0^+}\dfrac{\|e_n+ts\|_n-1}{t} = \max\{\mathrm{Re}(\lambda)\,\, |\,\, \lambda\in W_1(\phi_n(s)),\ \phi \in S_n(\mathcal S)\}.\]
Note that $\phi\in S_n(\mathcal S)$ has range in $M_n$, the elements $\lambda$ in the above equation is taken from the numerical range of the $n^2\times n^2$ matrix
$\phi_n(s_2)$.

In Section \ref{proofs}, we will provide the proofs of Theorem \ref{thm1} and Theorem \ref{cor3}. In Section \ref{section3}, applications of these results will be given in the commutative C$^*$-algebras and simple finite-dimensional C$^*$-algebras. A few new results, for instance Theorem \ref{thm3} and Theorem \ref{thm5}, that are useful in the geometry of Banach spaces, are also mentioned. 

\section{Proofs of the Main Results}\label{proofs}

\subsection{Proof of Theorem \ref{thm1}}

Our assumptions are that $V$ is an operator space with matrix norms $\|\cdot\|_n$, and that $n\in\mathbb N$, $v,w\in M_n(V)$ are fixed.
We aim to 
show the Gateaux derivative of $\|\cdot\|_n$ at $v$ in the direction of $w$, namely the quantity
\begin{equation}\label{e:gd1}
\lim_{t\rightarrow 0^+} \dfrac{\|v+tw\|_n-\|v\|_n}{t},
\end{equation}
 is given by
\begin{equation}\label{e:gd2}
\max\left\{\bra\Re(\phi_n(w))\eta, \eta\ket \, |\, \phi  \text{ is a support mapping for } v \text{ and }\eta\in M_0(\phi_n(v))\right\}.
\end{equation}

Because
$\|\cdot\|_n$ is a convex function, we have 
\[
\lim_{t\rightarrow 0^+}\dfrac{\|v+tw\|_n-\|v\|_n}{t}=\inf_{t\geq 0}\left\{\dfrac{\|v+tw\|_n-\|v\|_n}{t}\right\}.
\]
If $\phi : V\rightarrow M_n$ is a support mapping for $v$, then
	\begin{align*}
		\|v+tw\|_n&\geq \rho^*(\Re(\phi_n(v+tw)))\\
			&=\sup\{\langle\Re(\phi_n(v+tw))\eta, \eta\rangle \,\,|\,\, \eta\in\mathbb C^n, \|\eta\|=1\}\\
			&\geq \sup\{\bra\Re(\phi_n(v+tw))\eta, \eta\ket \,:\,\eta\in M_0(\phi_n(v))\}\\
			&=\sup\{\bra\Re(\phi_n(tw))\eta, \eta\ket \,:\,\eta\in M_0(\phi_n(v))\}+\|v\|_n,
	\end{align*}
which establishes the inequality
\[
\dfrac{\|v+tw\|_n-\|v\|_n}{t}\geq \sup\{\bra\Re(\phi_n(w))\eta, \eta\ket \,:\,\eta\in M_0(\phi_n(v))\}.
\]
By considering the infimum over all $t\geq 0$ and the supremum over all support functionals $\phi : V\rightarrow M_n$ of $v$, 
we deduce the quantity in (\ref{e:gd2}) is less than or equal to the quantity in (\ref{e:gd1}).

To prove the complementary inequality, 
let 
\[
a=\lim_{t\rightarrow 0^+}\dfrac{\|v+tw\|_n-\|v\|_n}{t}.
\]
By the Hahn-Banach Theorem and the convexity of the norm function,
there exists $F\in (M_n(V))^*$ such that 
\begin{equation}\label{step1}
\|F\|=1,\, \mathrm{Re}(F(v))=\|v\|_n,\text{ and } \mathrm{Re}(F(w))=a.
\end{equation} 
Arguing along the lines of the proof of \cite[Theorem 3.3.2]{winkler1996} (explained below), 
there are a support mapping $\phi : V\rightarrow M_n$ for $v$ and 
unit vector $\eta\in (\mathbb C^n)^n$ such that 
\begin{equation}\label{laststep}
F(u)=\bra \phi_n(u)\eta, \eta\ket, \text{ for all } u\in V.
\end{equation} 
Thus, 
\[
a=\mathrm{Re}(\bra\phi_n(w)\eta, \eta\ket)=\bra \Re(\phi_n(w))\eta, \eta\ket \leq \sup\{\bra\Re(\phi_n(w))\gamma, \gamma\ket \,:\,\gamma\in M_0(\phi_n(v))\},
\]
thereby showing 
the quantity in (\ref{e:gd1}) is less than or equal to the quantity in (\ref{e:gd2}), and establishing the assertion stated in Theorem \ref{thm1}.

Therefore, to complete the proof, we adapt the argument in \cite[Theorem 3.3.2]{winkler1996} to show
there are a support mapping $\phi : V\rightarrow M_n$ for $v$ and 
unit vector $\eta\in (\mathbb C^n)^n$ such that $F(u)=\bra \phi_n(u)\eta, \eta\ket$, for all $u\in V$.

Assume
$F\in (M_n(V))^*$ is a linear functional with the properties exhibited in \eqref{step1}.
Using \cite[Lemma 2.3.5]{winkler1996} (or drawing from the proof of \cite[Lemma 2.3.2]{Effros--Ruan-book}), there is a state
$p$ on $M_n$ such that 
\begin{equation}\label{step2}\mathrm{Re }F(A^*uA)\leq p(A^*A)\rho_k(u),
\end{equation} 
for all $u\in M_n(V)$ and $A\in M_{k,n}$, for arbitrary $r\in\mathbb N$. 
Since $p$ is a linear
functional on $M_n$, there exist $\eta_1,\dots,\eta_n\in\mathbb C^n$ such that 
\[
p(A)=\sum\limits_{i=1}^n \bra A\eta_i,\eta_i\ket,
\]
for all $A\in M_n$. Letting $\eta=(\eta_1,\dots,\eta_n)\in\mathbb C^n$, we get $$p(A)=\bra (A\tens{}I_n)\eta, \eta\ket.$$ Since $p$ is a state on $M_n$, this implies $\|\eta\|\leq 1$. Thus, $$p(A^*A)=\bra (A^*A\tens{}I_n)\eta,\eta\ket=\|(A\tens{}I_n)\eta\|^2.$$ Using \eqref{step2}, this implies \begin{equation}\label{step2a}|F(A^*uA)|\leq \sigma(A)\|(A\tens{}I_n)\eta\|^2,\end{equation} for all $A\in M_n$, where $\sigma(A)=2\max\{\rho_r(\pm A),\rho_r(\pm iA)\}$. Let $B_u: \mathcal H\rightarrow\mathcal H$ be the sesquilinear form defined on the subspace $\mathcal H=(M_n\tens{}I_n)\eta$ of $\mathbb C^{n^2}$ as $$B_u((B\tens{}I_n)\eta, (A\tens{}I_n)\eta)=F(A^* uB).$$ Then, $B_u$ is a bounded sesquilinear form  using \eqref{step2a}, and thus as an application of Riesz Representation Theorem there exists a bounded operator  $\psi(u)$ on $\mathcal H$ such that $$F(A^*uB)=\bra \psi(u)(B\tens{}I_n)\eta, (A\tens{}I_n)\eta\ket.$$

	Clearly, $\psi$ is a linear map. If we consider $P$ to be the projection from $\mathbb C^{n^2}$ onto $\mathcal H$, then $P\in(M_n\tens{}I_n)'$ (since $\mathcal H$ is invariant under $M_n\tens{}I_n$). Let $\tilde{\phi} : V\rightarrow M_n$ be the map defined by $$\tilde{\phi}(u)=P\psi(u)P,$$ and this gives $$F(A^*uB)=\bra \tilde{\phi}(u)(B\tens{}I_n)\eta, (A\tens{}I_n)\eta\ket=\bra \tilde{\phi}(u)\eta, \eta\ket,$$ for all $A, B\in M_n$. In particular, we have  $$F(u)=\bra \tilde{\phi}(u)\eta, \eta\ket,\  \tilde{\phi}(uA)=\tilde{\phi}(u)(A\tens{}I_n), \text{ and }\tilde{\phi}(A u)=(A\tens{}I_n)\tilde{\phi}(u),$$ for all $u\in M_n(V)$ and $A\in M_n$. Using linear algebra techniques, these in turn implies that there exists $\phi : V\rightarrow M_n$ such that $\tilde{\phi}(u)=\phi_n(u)$. Thus, we get $$F(A^*uB)=\bra \phi_n(u)\eta, \eta\ket,$$ for all $A,B\in M_n$, proving \eqref{step1} and, hence, the theorem.
\qed

\subsection{Proof of Theorem \ref{cor3}}

We aim to prove, for $s_1, s_2\in M_n(\mathcal S)$, where
$\mathcal S$ is an operator system, that
\begin{align*}
	\lim\limits_{t\rightarrow 0^+}\dfrac{\|s_1+ts_2\|_n-\|s_1\|_n}{t} = \frac{1}{\|s_1\|_n}\max\{& \mathrm{Re}(\bra\phi_n(s_2)\eta, \phi_n(s_1)\eta\ket) \,\, |\,\,  \phi \in S_n(\mathcal S)\\
	&\|\phi_n(s_1)\|=\|s_1\|_n\text{ and } \eta\in M_0(\phi_n(s_1))\}.
\end{align*}

Let $\phi\in S_k(\mathcal S)$ with $\|\phi_n(s_1)\|=\|s_1\|_n$, $A_1:=\phi(s_1)\in M_n(M_k(\mathcal S))$, and $A_2:=\phi_n(s_2)\in M_n(M_k(\mathcal S))$. Let $\eta\in M_0(A_1)$, i.e., $\bra\Re(A_1)\eta,\eta\ket=\|A_1\|$. As an application of Cauchy-Schwarz inequality, we have $$\|A_1\|=\bra\Re(A_1)\eta,\eta\ket=\mathrm{Re}(\bra A_1\eta,\eta\ket)\leq |\bra A_1\eta,\eta\ket|\leq \|A_1\eta\|\leq \|A_1\|.$$ So, equality occurs throughout and using the condition of equality in Cauchy-Schwarz inequality, we get $$\eta =\frac{1}{\|A_1\|}A_1\eta=\frac{1}{\|s_1\|_n}A_1\eta.$$

Further, we note that $\|\phi_n\|=1$.
Consequently, we  have 
\begin{align*}
	\|s_1+ts_2\|_n &\geq \|\phi_n(s_1+ts_2)\|\\
	&=\|A_1+tA_2\|\\
	&\geq \mathrm{Re}(\bra (A_1+tA_2)\eta, \eta\ket)\\
	&=\mathrm{Re}(\bra A_1\eta, \eta\ket)+t\mathrm{Re}(\bra A_2\eta, \eta\ket)\\
	&=\|s_1\|_n+t\frac{1}{\|s_1\|_n}\mathrm{Re}(\bra A_2\eta, A_1\eta\ket).
\end{align*}
This proves $\dfrac{1}{\|s_1\|_n}\mathrm{Re}(\bra A_2\eta, A_1\eta\ket) \leq \dfrac{\|s_1+ts_2\|_n-\|s_1\|_n}{t}$ for all $t>0$, which implies the quantity
on the right hand side of \eqref{e:qaz} is no larger than the quantity on the left hand side of \eqref{e:qaz}.

Using Theorem \ref{thm1}, there exist a support mapping
$\psi :\mathcal S\rightarrow M_n$
for $s_1$ and a vector $\eta\in M_0(\phi_n(e))$ such that
\[
\lim\limits_{t\rightarrow 0^+}\dfrac{\|s_1+ts_2\|_n-\|s_1\|_n}{t}=\bra \Re(\psi_n(s_2))\eta, \eta\ket\quad \text{and}\quad \bra \Re(\psi_n(s_1))\eta,\eta\ket=\|s_1\|.
\]
Let $\mathcal A$ be a $C^*$-algebra containing $\mathcal S$ is an operator subsystem. 
By Wittstock's extension theorem \cite[Theorem 8.2]{Paulsen-book}, we can extend $\psi$ to a completely contractive map to $\mathcal A$ that we will denote by same symbol. Using Wittstock's decomposition theorem \cite[Theorem 8.5]{Paulsen-book}, there exists a Hilbert space $\mathcal K$, a $\ast$-representation $\pi : \mathcal A\rightarrow\mathcal B(\mathcal K)$, and linear isometries $V_1, V_2:\mathbb C^n\rightarrow\mathcal K$ such that 
\[
\psi(a)=V_1^\ast\pi(a)V_2\text{ for all }a\in\mathcal A.
\]
 Thus, 
 \begin{equation}\label{eq1}\lim\limits_{t\rightarrow 0^+}\dfrac{\|s_1+ts_2\|_n-\|s_1\|_n}{t}=\mathrm{Re}(\bra \pi_n(s_2)\widetilde{V}_2\eta, \widetilde{V_1}\eta\ket)
 \end{equation} 
 and  $\mathrm{Re}(\bra \pi_n(s_1)\widetilde{V}_2\eta,\widetilde{V_1}\eta\ket)=\|s_1\|_n$ where $\widetilde{V}_i=V_i\oplus\dots\oplus V_i$ for $i=1,2$. We note that $\mathrm{Re}(\bra \pi_n(s_1)\widetilde{V}_2\eta,\widetilde{V_1}\eta\ket)=\|s_1\|$ implies $\bra \pi_n(s_1)\widetilde{V}_2\eta,\widetilde{V_1}\eta\ket=\|s_1\|_n$, since $\pi_n$ is contractive.
 
Using the condition of equality in Cauchy-Schwarz inequality and contractivity of $\pi_n$, the equation $\bra \pi_n(s_1)\widetilde{V}_2\eta,\widetilde{V_1}\eta\ket=\|s_1\|_n$ would imply 
 \begin{equation}\label{eq2}\widetilde{V}_1\eta=\frac{1}{\|s_1\|_n}\pi_n(s_1)\widetilde{V}_2\eta.
 \end{equation}
 
Now, we consider $\phi : \mathcal S\rightarrow M_n$ as 
\[
\phi(x)=V_2^*\pi(x)V_2\quad\text{ for all }\quad x\in\mathcal S.
\]
Then, $\phi$ is completely positive map. Since $V_2$ is an isometry and $\pi$ is $\ast$-homomorphism, 
we get $\phi\in S_n(\mathcal S)$. Furthermore, 
\[
\frac{1}{\|s_1\|_n}\bra \phi_n(s_2)\eta, \phi_n(s_1)\eta\ket=\frac{1}{\|s_1\|_n}\bra \pi_n(s_2)\widetilde{V}_2\eta, \pi_n(s_1)\widetilde{V}_2\eta\ket=\bra \pi_n(s_2)\widetilde{V}_2\eta, \widetilde{V}_1\eta\ket,
\]
where we have used \eqref{eq2} in the last equation. The above equation along with \eqref{eq1} yield
\[
\lim\limits_{t\rightarrow 0^+}\dfrac{\|s_1+ts_2\|_n-\|s_1\|_n}{t}=\frac{1}{\|s_1\|_n}\bra \phi_n(s_2)\eta, \phi_n(s_1)\eta\ket.
\]
Now, $\phi\in S_n(\mathcal S)$ and $$\bra \phi_n(s_1)\eta, \phi_n(s_1)\eta\ket =\bra \pi_n(s_1)\widetilde{V}_2\eta, \pi_n(s_1)\widetilde{V}_2\eta\ket=\|s_1\|\bra \pi_n(s_1)\widetilde{V}_2\eta, \widetilde{V}_1\eta\ket=\|s_1\|^2.$$ This proves $\|\phi_n(s_1)\eta\|=\|s_1\|$. Since $\|\phi_n\|\leq 1$ and $\|\eta\|=1$, we get $\|\phi_n\|=\|s_1\|_n$. This completes the proof.
 \qed

\section{Applications of the Main Results}
\label{section3}

As an application of Theorem \ref{cor3}, we get the following special case of \cite[Theorem 2.4]{keckic2005}. Let $\|\cdot\|$ denotes the operator norm on $M_n$.

\begin{corollary}\label{cor1}
	Let $A, B\in M_n$. Then, $$\lim_{t\rightarrow 0^+}\dfrac{\|A+tB\|-\|A\|}{t}=\frac{1}{\|A\|}\max\{\mathrm{Re}(\bra B\eta, A\eta\ket) \, |\, \eta\in\mathbb C^n,\   \|\eta\|=1  \text{ and } \|A\eta\|=\|A\|\}.$$
\end{corollary}
\begin{proof} 
	It is easy to see that the right hand side of the equation is less than equal to the left hand side. To prove the other side inequality, we use Theorem \ref{cor3} with $V=\mathbb C$. Note that if $\phi\in S_n(\mathbb C)$, then $\phi(\lambda)=\lambda I_n$ for all $\lambda\in\mathbb C$ and $\phi_n(C)=C\tens{}I_n$ for all $C\in M_n$. Now, let $\phi\in S_n(\mathbb C)$ with $\|\phi_n(A)\|=\|A\|$ and $\gamma\in M_0(\phi_n(A))\subseteq (\mathbb C^n)^n$ such that $$\lim_{t\rightarrow 0^+}\dfrac{\|A+tB\|-\|A\|}{t}=\frac{1}{\|A\|}\mathrm{Re}(\bra (B\tens{}I_n)\gamma, (A\tens{}I_n)\gamma\ket).$$ Then we can decompose $\gamma$ as $\eta_1,\dots,\eta_n\in\mathbb C^n$ such that $\sum_{i=1}^n\|\eta_i\|^2=1$, $$\bra (B\tens{}I_n)\gamma, (A\tens{}I_n)\gamma\ket=\sum_{i=1}^n \bra B\eta_i, A\eta_i\ket\quad\text{and}\quad \|A\|=\bra \Re(A\tens{}I_n)\gamma, \gamma\ket=\sum_{i=1}^n\bra \Re(A)\eta_i, \eta_i\ket.$$ Using Cauchy-Schwarz inequality we have $$\|A\|=\sum_{i=1}^n\|\eta_i\|^2\left\bra \Re(A)\frac{\eta_i}{\|\eta_i\|}, \frac{\eta_i}{\|\eta_i\|}\right\ket\leq \sum_{i=1}^n\|A\eta_i\|\|\eta_i\|^2\leq  \|A\|\sum_{i=1}^n\|\eta_i\|^2=\|A\|,$$ which using the condition of equality in   Cauchy-Schwarz inequality implies $\|A\eta_i\|=\|A\|\|\eta_i\|$ whenever $\eta_i\neq 0$. Thus, $$\bra (B\tens{}I_n)\gamma, (A\tens{}I_n)\gamma\ket=\sum_{\eta_i\neq 0} \bra B\eta_i, A\eta_i\ket=\sum_{\eta_i\neq 0}\|\eta_i\|^2\left\bra B\frac{\eta_i}{\|\eta_i\|}, A\frac{\eta_i}{\|\eta_i\|}\right\ket,$$ lies in the convex hull of numerical range of in the numerical range of $(A^\ast B)\tens{} I_n$ restricted to the subspace $\{\eta\in \mathbb C^n : \|A\eta\|=\|A\| \}$. Using Haudorff-Toeplitz theorem, there exists $\eta\in\mathbb C^n$ such that $\bra (B\tens{}I_n)\gamma, (A\tens{}I_n)\gamma\ket=\bra B\eta, A\eta\ket$. Thus, $$\lim_{t\rightarrow 0^+}\dfrac{\|A+tB\|-\|A\|}{t}=\frac{1}{\|A\|}\mathrm{Re}(\bra (B\tens{}I_n)\gamma, (A\tens{}I_n)\gamma\ket)=\frac{1}{\|A\|}=\bra B\eta, A\eta\ket.$$ This completes the proof.
\end{proof}

Gateaux derivatives are closely associated with a notion of orthogonality in a normed space, known as Birkhoff-James \cite{james1947}. Let $\mathcal X$ be a vector space. Let $x\in \mathcal X$ and $\mathcal Y$ be a subspace of $X$. Then, $x$ is said to be (Birkhoff-James) orthogonal to $\mathcal Y$ if and only if $$\|x\|\leq \|x-y\|\quad\text{ for all }\quad y\in \mathcal Y.$$ When $\mathcal Y=\mathbb Cy$ is one dimensional subspace generated by $y\in \mathcal X$, $x$ is orthogonal to $\mathcal Y$ is abbreviated as $x$ is orthogonal to $y$. It was proved in \cite[Proposition 1.5]{keckic2005} that $x$ is orthogonal to $\mathbb Cy$ if and only if $\inf_{\varphi\in[0,2\pi]}D_{\varphi, x}(y)\geq 0$, where $$D_{\varphi, x}(y)=\lim_{t\rightarrow 0^+}\dfrac{\|x+te^{i\varphi}y\|-\|x\|}{t}.$$ Using this fact, Corollary \ref{cor1}, and along the lines of \cite[Corollary 3.1]{keckic2005}, we get the following.

\begin{corollary}\label{BS}
	Let $A\in M_n$, then $A$ is orthogonal to $B$ if and only if there exists $\eta\in \mathbb C^n$ such that $\|\eta\|=1$, $\bra A\eta, \eta\ket=\|A\|$ and $\bra B\eta, \eta\ket =0$.
\end{corollary}

\begin{remark} This theorem has become famous by the name of Bhatia-\v Semrl Theorem after this appeared in \cite[Theorem 1.1]{bhatia--semrl1999} in the following form: For $B\in M_n$, we have $\|A\|\leq \|A+\lambda B\|$ for all $\lambda\in\mathbb C$ if and only if there exists $\eta\in \mathbb C^n$ such that $\|\eta\|=1$, $\bra A\eta, \eta\ket=\|A\|$ and $\bra B\eta, \eta\ket =0$.
	
This theorem is actually a special case of \cite[Lemma 2.2]{magajna1993}, that was motivated by the proof of \cite[Theorem 2]{stampfli1970} and \cite[Lemma 9.14]{Davidson-book-nestalgebras}.
\end{remark}

Now, we see a generalization of this form of Bhatia-\v Semrl Theorem in the setting of operator spaces.
\begin{theorem}\label{thm3}
	Let $(V, \|\cdot\|_n)$ be an operator space. Let $v\in M_n(V)$ and $\mathcal W$ be a subspace of $M_n(V)$. Then, we have $$\|v\|_n\leq \|v-w\|_n\quad\text{ for all }\quad w\in\mathcal W$$ if and only if there is a completely contractive mapping $\phi : V\rightarrow M_n$ and a unit vector $\eta\in(\mathbb C^n)^n$ such that $$\|\phi_n(v)\eta\|=\|v\|_n\quad\text{and}\quad \bra\phi_n(v)\eta, \phi_n(w)\eta\ket=0\quad\text{for all }\quad w\in\mathcal W.$$
\end{theorem}
\begin{proof} The converse is easy. If $\phi : V\rightarrow M_n$ is a completely contractive mapping and  $\eta\in(\mathbb C^n)^n$ is a unit vector such that $$\|\phi_n(v)\eta\|=\|v\|_n\quad\text{and}\quad \bra\phi_n(v)\eta, \phi_n(w)\eta\ket=0\quad\text{for all }\quad w\in\mathcal W.$$ Then, \begin{align*}
		\|v\|_n^2&=\bra\phi_n(v)\eta, \phi_n(v)\eta\ket\\
		&\leq \bra\phi_n(v)\eta, \phi_n(v)\eta\ket+\bra\phi_n(w)\eta, \phi_n(w)\eta\ket\\
		&=\bra\phi_n(v+w)\eta, \phi_n(v+w)\eta\ket\\
		&\leq \|v+w\|_n^2.
	\end{align*}
	Now, we prove the other direction. Let $\|v\|_n\leq \|v-w\|_n$ for all $w\in\mathcal W$. 
		Using \cite[Theorem 2.1]{james1947},  there exists a functional $f\in M_n(V)^*$ such that $\|f\|=1$, $f(v)=\|v\|$ and $f(w)=0$ for all $w\in \mathcal W$. Using \cite[Lemma 2.3.3]{Effros--Ruan-book}, there exists a complete contraction $\phi : V\rightarrow M_n$ and unit vectors $\xi,\eta\in(\mathbb C^n)^n$ such that \begin{equation}\label{equation1} f(u)=\bra\phi_n(u)\eta, \xi\ket,\end{equation} for all $u\in M_n(V)$. Now, a simple application of the condition of equality in Cauchy-Schwarz inequality implies that for a functional of the form \eqref{equation1}, the condition  $f(v)=\|v\|_n$ implies $\xi=\frac{1}{\|v\|_n}\phi_n(v)\eta.$ Combining all these statements give the required result.
\end{proof}

In case of operator system, we get the following result.

\begin{corollary}\label{thm4}
	Let $\mathcal S$ be an operator system. Let $s\in\mathcal S$ and $\mathcal W$ be a subspace of $M_n(\mathcal S)$. Then, we have $$\|s\|_n\leq \|s-w\|_n\quad\text{ for all }\quad w\in \mathcal W$$ if and only if there exists a matrix state $\phi\in S_n(\mathcal S)$ 
	and a unit vector $\eta\in(\mathbb C^n)^n$ such that $$\|\phi_n(s)\eta\|=\|s\|_n\quad\text{and}\quad \bra\phi_n(s)\eta, \phi_n(w)\eta\ket=0\quad\text{for all }\quad w\in\mathcal W.$$
\end{corollary}
\begin{proof}
	This follows along the lines of proof of Theorem \ref{cor3} and using Theorem \ref{thm3}.
\end{proof}

The above corollary can be used to the existence of some special states in operator systems and $C^*$-algebras. We recall \cite{Rockafellar-book} that a norm's face $F$ of $\mathcal S$  is a convex subset of the unit ball of $\mathcal S$ such that if for $s_1,s_2\in\mathcal S$ with $\|s_1\|=\|s_2\|=1$ and $\lambda s_1+(1-\lambda)s_2\in F$ with $0<\lambda<1$, then $s_1,s_2\in F$. A norm's face is maximal if it is not properly contained in any other norm's face. Let $\mathrm{ri}(\mathcal F)$ denotes the relative interior of a norm's face.

\begin{corollary}\label{cor6}
		Let $\mathcal S$ be an operator system with Archimedean order unit $e$. Let $\mathcal F$ be the maximal norm's face of $\mathcal S$ containing $e$. If $\mathcal W$ is a subspace of $\mathcal S$ such that 		
		 $$1\leq \|e- w\|\quad\text{ for all }\quad w\in \mathcal W,$$  then there exists a matrix state $\phi\in S_1(\mathcal S)$ such that $\phi(v)=\|v\|$ for all $v\in\mathrm{ri}(\mathcal F)$ and $0\in W_1(\phi(w))$ for all $w\in\mathcal W$.
\end{corollary} 
\begin{proof}
	Using Corollary \ref{thm4}, there exists a state $\phi\in S_1(\mathcal S)$ such that $0\in W_1(\phi(w))$ for all $w\in\mathcal W$. We have, $\phi(e)=1$. If we prove $\phi(v)=\|v\|$ for all $v\in\mathrm{ri}(\mathcal F)$ and we will be done. Let $v\in\mathrm{ri}(\mathcal F)$. Using \cite[Theorem 6.2]{Rockafellar-book},  there exists $\epsilon>0$ such that $$\|(1-t)e+tv\|=1\text{ for all }t\in [0,1+\epsilon].$$ Then, $$|t\phi(v)+(1-t)|=|\phi((1-t)e+t s)|\leq \|(1-t)e+t v\|=1,$$ for all $t\in [1-\epsilon,1+\epsilon]$. Taking $t=1\pm \epsilon$, we get $|1\pm \epsilon(f(v)-1)|\leq 1$, which gives $f(v)=1$.
\end{proof}

For a $C^*$-algebra $\mathcal A$, a slightly different version of the above corollary also holds.

\begin{theorem}\label{thm5}
		Let $\mathcal A$ be a unital $C^*$-algebra with unit element $1_{\mathcal A}$. Let $x\in\mathcal A$ be an element of norm one and $\mathcal F$ be the maximal norm's face of $\mathcal A$ containing $x$. If $\mathcal B$ is a subspace of $\mathcal A$ such that $$\|x\|\leq \|x-y\|\quad\text{ for all }\quad y\in \mathcal B,$$  then there exists a state $\phi\in S_1(\mathcal A)$ such that $$\phi(x^*x)=1\text{ and } \phi(x^*y)=0\quad \text{ for all }\quad x\in\mathrm{ri}(\mathcal F)\text{ and }y\in \mathcal B.$$
\end{theorem}
\begin{proof}
	By Hahn-Banach theorem, there exists a functional $f\in\mathcal A^*$ such that $\|f\|=1=f(x)$ and $f(y)=0$ for all $y\in\mathcal B$. Using the argument in Corollary \ref{cor6}, we have $f(x)=1$ for all $x\in\mathrm{ri}(\mathcal F)$. For functional $f$, there exists a cyclic representation $(\mathcal H, \pi, \xi)$ with $\|\xi\|=1$ and a unit vector $\eta\in\mathcal H$ such that $$f(z)=\bra\pi(z)\xi, \eta\ket\quad\text{ for all }z\in\mathcal A.$$ Since $f(x)=1$ , we get $\bra\pi(x)\xi, \eta\ket=\|x\|$ for all $x\in\mathrm{ri}(\mathcal F)$. Using the condition of Cauchy-Schwarz inequality,
	we get $\eta=\pi(x)\xi$ for all $x\in\mathrm{ri}(\mathcal F)$. If we take $$\phi(z)=\bra \pi(z)\xi,\xi\ket\quad\text{ for all}\quad z\in\mathcal A,$$ then $\phi$ is the required state on $\mathcal A$.
\end{proof}

Finally, we see some matricial versions of some known results. We recall the notion of POVMs (see \cite{farenick--plosker--smith2011} for more details). Assume that $X$ is a compact Hausdorff space, and $\mathcal O(X)$ is the $\sigma$-algebra of Borel sets of $X$. Let $\mathcal B(\mathcal H)$ denotes the space of bounded linear operators on a Hilbert space $\mathcal H$. A function $\nu : \mathcal O(X)\rightarrow \mathcal B(\mathcal H)$ is a POVM on $X$ if \begin{enumerate}
	\item[(i)] $\nu(E)\in\mathcal B(\mathcal H)_+$ with $0\leq \lambda\leq 1$ for every eigenvalue $\lambda$ of $\nu(E)$;
	\item[(ii)] for every countable collection $\{E_k\}_{k\in\mathbb N}\subseteq\mathcal O(X)$ with $E_k\cap E_{k'}=\emptyset$ for $k\neq k'$, $$\nu\bigg(\cup_{n\in\mathbb N}E_k\bigg)=\sum_{k\in\mathbb N}\nu(E_k),$$ where the convergence 
	of the right-hand side of the equality above is with respect to the weak
	operator topology of $\mathcal B(\mathcal H)$.
\end{enumerate}
If a POVM $\nu$ on $\mathcal O(X)$ satisfies $\nu(X)=I_{\mathcal B(\mathcal H)}$, then it is referred to as a \emph{quantum probability measure}. If $\nu:\mathcal O(X)\rightarrow M_n$ is a positive operator-valued measure (POVM), then, for each pair $\xi,\nu\in\mathbb C^n$, the formula
\[
\mu_{\xi,\eta}^{[\nu]}(E)=\langle \nu(E)\xi,\eta\rangle, \mbox{ for }E\in\mathcal O(X),
\]
defines a complex Borel measure on $\mathcal O(X)$. Hence, for each $f\in C(X)$, we denote by $\displaystyle\int_Xf\,d\nu$
the unique operator on $\mathbb C^n$ for which
\[
\left\langle \left(\displaystyle\int_Xf\,d\nu\right)\xi,\eta\right\rangle= \int_X f\,d\mu_{\xi,\eta}^{[\nu]}, \mbox{ for all }\xi,\eta\in\mathbb C^n.
\] 

We get the following extension of \cite[Theorem 1.3]{Singer-book} (see also \cite[Corollary 2.1]{keckic2012}).

\begin{corollary}\label{doug}
	If $X$ is a compact Hausdorff space and $Id_n$ denotes the diagonal matrix in $M_n(C(X))$ with diagonal entries $1_X$. The, for a subspace $\mathcal W$  of  $M_n(C(X))$, we have
\[ \|Id_n\|_{M_n(C(X))}\leq \|Id_n- G\|_{M_n(C(X))}, \mbox{ for all }G\in \mathcal W,
\]
if and only if there exists a quantum probability measure 
$\nu : \mathcal O(X)\rightarrow M_n\otimes M_n$ and a unit vector $\eta\in\mathbb C^n\otimes\mathbb C^n$ such that
$\left\langle \left(\displaystyle\int_X G\,d\nu\right)\eta,\eta\right\rangle=0$, for all $G\in\mathcal W$.
\end{corollary}
\begin{proof}
	Assume $\nu : \mathcal O(X)\rightarrow M_n\otimes M_n$ is a quantum probability and $\eta\in\mathbb C^n\otimes \mathbb C^n$ 
	is a unit vector
	for which $\langle (\int_X G\,d\nu)\eta,\eta\rangle=0$, for all $G\in\mathcal W$. 
	If $\{\epsilon_i\}_{i=1}^n$ denotes the canonical orthonormal basis for $\mathbb C^n$, then there exist $\eta_1,\dots,\eta_n\in\mathbb C^n$
	such that
	\[
	\eta = \sum_{i=1}^n \eta_i\otimes \epsilon_i.
	\]
	Let $W:\mathbb C^n\rightarrow \mathbb C^n \otimes \mathbb C^n$ be the linear map for which $We_i=\eta_i\otimes \epsilon_i$, for each $i$.
	Let $\xi\in\mathbb C^n$ denote the unique unit
	vector for which $W\xi=\eta$. The map $W^*\nu W:\mathcal O(X)\rightarrow M_n$ is a POVM and has the property that, for every $E\in\mathcal O(X)$
	and $\gamma,\delta\in\mathbb C^n$,
	\[
	\mu_{\gamma,\delta}^{[W^*\nu W]}(E)=\langle W^*\nu(E) W\gamma,\delta \rangle = \mu_{W\gamma,W\delta}^{[\nu]}(E).
	\]
	In particular,
	\[
	\mu_{\xi,\xi}^{[W^*\nu W]}(E) = \mu_{\eta,\eta}^{[\nu]}(E).
	\]
	Now define a linear map $\phi:C(X)\rightarrow M_n$ by
	\[
	\phi(f)=\int_X f\,d(W^*\nu W).
	\]
	Note that $\phi$ is a matrix state. If $F=[f_{ij}]_{i,j}\in C(X)\otimes M_n$, then
	\[
	\begin{array}{rcl}
	\langle \phi_n(F)\eta,\eta\rangle &=& \displaystyle\sum_{i,j=1}^n \langle \phi(f_{ij})\eta_j,\eta_i\rangle \\ && \\
	&=& \displaystyle\sum_{i,j=1}^n \displaystyle\int_X f_{ij} \, d\mu_{W\epsilon_j,W\epsilon_i}^{[\nu]} \\ && \\
	&=& \left\langle\left( \displaystyle\int_X F \, d\nu\right)\eta,\eta \right\rangle.	
	\end{array}
	\]
	Hence, $\langle\phi_n(G)\eta,\eta\rangle=0$, for every $G\in\mathcal W$, and so, by Corollary \ref{thm4},
	$$\|Id_n\|_{M_n(C(X))}\leq \|Id_n- G\|_{M_n(C(X))}, \mbox{for every }G\in\mathcal W.$$
	
	Conversely, assume that $\|Id_n\|_{M_n(C(X))}\leq \|Id_n- G\|_{M_n(C(X))}$, for every $G\in\mathcal W$. By Corollary \ref{thm4}, there exist a 
	matrix state $\phi:C(X)\rightarrow M_n$ and a unit vector $\eta\in\mathbb C^n\otimes\mathbb C^n$ such that
	$\langle\phi_n(G)\eta,\eta\rangle=0$, for all $G\in\mathcal W$. As noted in \cite[\S4]{Paulsen-book}, the matrix state $\phi$ is 
	determined by a unique quantum probability measure
	$\omega:\mathcal O(X)\rightarrow M_n$ whereby, for each pair $\gamma,\delta\in \mathbb C^n$,
	\[
	\langle\phi(f)\gamma,\delta\rangle =\int_X f\,d\mu_{\gamma,\delta}^{[\omega]},
	\mbox{ for all } f\in C(X).
	\]
	(The POVM $\omega$ has the form $\omega(E)=V^* P(E) V$, where $V^*\pi V$ is a minimal
	Stinespring dilation of $\phi$ and $P$ is the spectral measure on $\mathcal O(X)$ arising from the Spectral Theorem applied to the abelian C$^*$-subalgebra
	$\pi\left(C(X)\right)$ of $\mathcal B(\mathcal H_\pi)$.) Define an operator-valued measure $\nu:\mathcal O(X)\rightarrow M_n\otimes M_n$ by
	\[
	\nu(E)=\sum_{i=1}^n\sum_{j=1}^n \omega(E)\otimes E_{ij},
	\]
	where $\{E_{ij}\}_{i,j=1}^n$ are the canonical matrix units of $M_n$. Note that $\nu$ is positive, as $\omega$ is positive; hence 
	$\nu$ is a quantum probability measure.
	
	Express $\phi_n$ as $\phi_n=\phi\otimes\mbox{\rm Id}_{M_n}$, so that, for every $F=(f_{ij})\in C(X)\otimes M_n$, we have
	\[
	\phi_n(F)=\sum_{i,j=1}^n \left(\int_X f_{ij}\,d\,\omega\right)\otimes E_{ij} = 	\int_X F\,d\,\nu.
	\]
	Because $\langle\phi_n(G)\eta,\eta\rangle=0$ if $G\in \mathcal W$, we deduce 
	$\left\langle \left(\displaystyle\int_X G\,d\nu\right)\eta,\eta\right\rangle=0$, for all $G\in\mathcal W$.
\end{proof}

Along the lines of proof of Corollary \ref{doug} and Theorem \ref{cor3}, we also get the following extension of \cite[Theorem 2.1]{keckic2012} for the Gateaux derivative in case of commutative $C^*$-algebras.

\begin{corollary}
	Let $X$ be a compact Hausdorff space, and $C(X)$ be the space of all continuous function on $X$. Let $F=(f_{ij}), G=(g_{ij})\in M_n(C(X))$, then \begin{align*}\lim\limits_{t\rightarrow 0^+}\dfrac{\|F+tG\|_n-\|F\|_n}{t}=\max\bigg\{&\mathrm{Re}\left\langle \left(\displaystyle\int_X G\,d\nu\right)\eta,\left(\displaystyle\int_X F\,d\nu\right)\eta\right\rangle\,\, | \,\, \nu \text{ is a quantum}\\
		&\text{probability measure with } \left\|\displaystyle\int_X F\,d\nu\right\|=\|F\|_{M_n(C(X))}\bigg\}.\end{align*}
\end{corollary}

We end our article with a remark that there are notions of real operator spaces and real operator systems, and all our results also extend naturally when the underlying field is $\mathbb R$. 

\subsection*{Acknowledgements}

The author would like to thank Douglas Farenick for many useful discussions. The author is also thankful for the great hospitality service provided by BIRS (Banff, Canada) during the workshop `Operator Systems and their Applications (25w5405)', where the idea of the project originated. 

 

\end{document}